x - 5/12 = 15/4 what does x equal?
Answer by stanbon(75887)

x - 5/12 = 15/4 what does x equal?
-----
x = (15/4) + (5/12)
x = (45/12) + (5/12)
x = 50/12
x = 25/6
==============
Cheers,
Stan H.